# Climb: Class-imbalanced Learning Benchmark on Tabular Data

### v1

@@ A.3 Additional Results
RUS CC IHT NM TL ENN RENN AKNN OSS NCR ROS SMT BSMT SSMT ASYN SPE BC BRF EE UBS UBA OBS SMBS OBA SMBA CS AdaC AdaBS AsyBS
bwin_amlb 36.4 34.4 35.9 36.9 35.6 35.3 39.6 39.6 36.4 35.4 38.6 39.5 36.3 37.1 35.4 35.9 38.6 37.2 40.7 39.0 38.0 39.0 38.3 38.7 36.6 37.6 34.9 36.2 34.7 36.2
mozilla4 82.9 78.3 76.7 65.6 66.9 82.0 81.7 80.9 80.1 81.9 81.7 83.4 81.8 82.1 82.1 79.5 90.3 89.5 88.5 89.5 89.6 89.5 84.1 81.7 90.4 89.7 83.4 83.7 83.8 83.7
mc2 40.7 43.9 38.4 39.8 43.9 43.7 46.9 45.9 41.2 43.7 45.9 42.1 45.9 43.8 50.4 50.6 50.9 46.4 47.0 51.8 51.1 51.8 54.2 55.1 51.9 52.2 41.1 48.2 46.7 48.2
wholesale-customers 72.5 70.0 68.0 60.2 68.1 71.6 76.5 76.6 75.4 72.2 76.5 72.9 74.0 73.3 77.6 71.6 83.6 81.3 81.6 82.6 81.7 82.6 71.6 77.8 83.1 82.3 74.4 74.2 75.9 74.2
vertebra-column 57.2 55.7 59.7 56.9 54.2 53.3 62.6 61.3 62.4 53.5 61.7 55.1 58.3 61.1 64.7 60.6 66.1 63.5 64.5 64.9 67.4 64.9 61.5 64.5 65.4 65.6 56.5 62.0 55.8 62.0
law-school-admission-bianry 37.4 37.5 34.3 40.0 37.0 38.2 39.5 40.4 40.2 38.2 40.0 37.1 37.6 37.6 37.7 37.7 42.7 41.6 43.1 44.1 42.9 44.1 39.2 39.4 42.7 42.1 37.4 39.4 39.0 39.4
bank32nh 46.4 45.4 44.0 47.0 45.7 45.7 47.9 48.0 48.2 46.7 47.8 45.4 45.7 46.4 47.4 46.0 57.2 57.4 55.7 56.6 56.4 56.6 46.8 45.7 54.9 56.4 45.7 47.2 46.3 47.2
elevators 56.9 52.0 46.0 47.4 45.4 56.2 56.6 56.5 55.3 56.4 56.6 55.6 56.6 56.6 56.2 56.8 71.1 70.3 65.8 67.4 68.3 67.4 56.6 57.7 69.6 70.0 55.9 57.2 56.0 57.2
cpu_small 70.9 69.1 69.5 57.8 61.5 72.1 73.3 72.7 72.5 72.7 72.9 72.0 71.7 70.4 71.3 69.5 80.6 79.4 78.3 79.0 80.0 79.0 71.4 72.1 80.2 79.9 71.0 71.6 71.3 71.6
Credit_Approval_Classification 38.1 36.3 36.6 39.4 38.6 41.2 41.2 39.2 39.7 41.2 40.3 37.6 38.8 38.1 39.1 40.0 46.6 44.7 46.1 47.5 46.0 47.5 39.0 38.7 45.4 46.4 39.1 39.0 40.6 39.0
house_8L 61.2 58.0 54.2 52.0 47.7 61.7 62.6 62.4 62.5 61.9 62.2 61.7 61.2 61.4 61.0 59.4 72.0 70.9 68.6 69.9 70.3 69.9 61.4 61.1 71.6 71.2 61.2 61.7 61.7 61.7
house_16H 61.3 58.0 55.6 51.6 46.7 61.6 62.1 62.3 62.2 61.6 62.1 60.1 60.2 61.3 59.9 58.8 72.5 71.6 69.3 70.6 70.7 70.6 59.2 60.7 72.1 71.4 59.3 61.6 59.1 61.6
phoneme 67.5 61.1 62.4 50.8 54.8 66.1 66.1 65.9 66.1 66.2 66.1 66.2 66.2 67.5 66.7 65.9 76.9 74.8 70.7 72.3 73.9 72.3 65.5 66.5 74.6 75.3 65.7 68.7 66.0 68.7
ilpd-numeric 33.8 35.4 38.6 38.4 30.8 33.6 40.2 39.6 39.4 33.8 41.4 38.4 38.3 37.2 34.8 35.8 43.6 38.2 45.0 44.0 43.8 44.0 33.5 37.2 35.0 39.2 38.2 32.6 36.5 32.6
planning-relax 27.7 31.7 27.4 28.5 30.9 31.0 29.9 30.3 29.8 30.1 31.4 27.6 33.0 31.1 30.7 33.3 28.8 28.1 28.5 30.1 34.0 30.1 29.4 32.3 31.9 32.1 30.7 29.4 30.4 29.4
MiniBooNE 71.0 69.2 65.0 58.2 61.3 71.7 70.4 68.5 69.5 71.6 70.9 70.6 70.5 69.3 70.3 69.6 82.6 81.5 78.9 80.1 77.6 80.1 77.7 77.5 81.0 81.3 70.6 77.5 77.6 77.5
machine_cpu 78.9 73.2 76.3 63.4 78.2 77.5 76.2 75.4 76.2 77.5 77.7 77.7 82.9 83.8 82.9 81.2 84.5 80.5 81.3 85.4 85.7 85.4 78.4 80.8 81.0 85.0 77.9 78.9 81.7 78.9
telco-customer-churn 36.3 38.0 36.0 39.3 29.0 38.4 41.8 43.3 42.4 38.5 41.5 37.5 38.2 38.9 38.6 39.0 45.7 44.7 46.0 46.7 46.0 46.7 40.9 40.0 45.3 45.9 37.5 37.9 39.0 37.9
haberman 30.3 28.1 27.6 32.6 32.2 34.4 36.5 34.8 34.2 34.7 36.5 30.6 32.2 32.2 30.3 30.6 33.6 31.9 35.1 36.0 32.7 36.0 32.3 34.2 33.1 36.4 31.0 30.5 31.7 30.5
vehicle 88.6 80.2 81.8 69.3 80.0 86.8 89.9 88.7 87.3 86.5 88.8 85.6 84.9 88.6 88.0 89.0 95.4 91.4 92.4 95.5 95.7 95.5 86.6 87.5 96.3 96.5 87.2 89.4 86.8 89.4
cpu 93.8 90.1 90.6 72.0 93.8 93.5 93.8 93.8 90.9 93.5 93.8 91.4 91.2 92.4 91.4 95.4 98.2 95.3 92.9 92.9 96.7 92.9 92.5 94.0 95.5 95.5 91.1 93.8 92.5 93.8
ada 45.0 45.1 43.4 41.1 36.1 46.9 50.6 51.5 50.4 47.3 49.0 46.5 46.8 45.1 45.4 44.8 57.4 56.8 54.6 57.1 55.4 57.1 46.8 46.8 58.0 57.7 45.6 45.9 46.3 45.9
adult 47.5 45.4 41.0 41.9 32.3 48.6 48.9 48.4 48.8 49.0 48.9 47.2 46.9 45.7 46.9 46.1 55.6 56.0 53.2 54.9 54.4 54.9 50.5 49.1 56.0 55.6 47.5 49.4 49.8 49.4
blood-transfusion-service-center 27.8 30.1 27.7 31.1 25.7 29.2 32.6 34.0 31.2 29.0 33.4 27.8 28.1 27.1 29.5 27.7 34.0 30.9 33.6 35.7 34.9 35.7 31.2 29.8 33.9 34.6 28.7 29.2 29.6 29.2
default-of-credit-card-clients 29.1 28.1 26.0 29.3 24.6 29.7 32.0 31.5 32.1 29.9 31.7 29.1 28.2 29.2 28.6 28.3 37.6 38.3 36.6 38.1 37.2 38.1 32.9 31.2 37.3 38.0 28.8 32.1 32.2 32.1
Customer_Churn_Classification 37.7 35.5 26.0 32.9 27.3 38.8 41.1 40.4 41.1 38.7 41.2 37.3 37.7 36.9 37.8 36.6 48.2 49.3 45.7 47.6 44.3 47.6 43.9 43.3 47.7 48.6 37.7 42.6 43.1 42.6
SPECTF 26.7 28.9 34.0 31.0 23.2 27.9 34.4 35.9 39.1 30.3 38.3 29.7 26.9 32.0 30.3 31.1 39.4 39.1 42.2 43.3 39.3 43.3 31.4 36.2 40.2 36.8 30.1 25.6 29.4 25.6
Medical-Appointment-No-Shows 25.0 25.6 20.1 27.2 23.2 25.1 26.4 27.3 26.8 25.1 26.3 24.9 24.3 24.6 24.6 24.3 28.2 26.8 28.8 28.3 28.0 28.3 24.7 24.4 25.0 25.8 24.8 25.0 24.8 25.0
JapaneseVowels 85.6 75.4 76.6 64.6 66.3 85.6 85.9 86.1 85.8 85.5 85.6 85.4 85.9 85.6 84.1 83.9 94.5 91.8 88.5 91.7 93.9 91.7 86.5 86.0 92.4 93.3 86.3 86.2 86.5 86.2
ibm-employee-attrition 24.2 21.1 18.8 23.8 18.2 22.2 25.1 25.1 24.8 22.4 24.8 21.8 25.8 25.3 24.6 24.4 36.6 33.9 32.3 34.9 34.4 34.9 22.8 24.6 26.1 29.9 20.7 23.5 22.4 23.5
first-order-theorem-proving 30.8 26.3 24.2 26.4 22.4 29.6 28.1 28.1 27.9 29.7 28.1 29.8 30.5 30.8 30.9 30.2 51.3 48.6 51.2 51.7 51.8 51.7 51.2 50.6 53.0 52.9 29.9 51.2 51.0 51.2
user-knowledge 71.1 63.4 62.6 54.8 60.1 67.5 69.0 69.0 67.2 67.5 69.0 68.5 74.2 71.1 67.3 71.1 89.4 89.0 86.0 86.2 90.2 86.2 72.4 73.9 86.7 86.7 69.2 74.0 74.2 74.0
online-shoppers-intention 38.1 36.7 24.6 31.3 25.2 38.9 41.8 40.9 41.3 39.1 41.3 38.6 38.7 37.9 40.0 39.2 50.0 50.4 46.2 48.1 47.7 48.1 37.3 39.9 47.5 50.5 37.7 37.2 37.6 37.2
kc1 24.2 24.4 14.9 23.9 20.7 24.3 25.9 27.5 23.7 24.5 26.7 23.4 27.8 26.9 29.0 26.6 31.0 26.8 28.3 30.2 30.8 30.2 29.4 28.6 28.3 30.6 22.1 25.2 24.9 25.2
thoracic-surgery 17.1 14.9 16.8 17.6 17.8 17.7 17.8 18.1 17.4 17.6 20.4 18.0 19.0 19.2 20.3 19.7 20.5 20.7 20.3 20.0 22.5 20.0 16.7 20.6 16.1 16.6 16.9 16.9 17.4 16.9
UCI_churn 53.4 35.3 21.4 27.2 31.7 52.2 54.4 54.5 54.4 52.6 54.2 54.2 53.0 53.5 52.2 51.4 74.6 71.0 58.3 65.9 65.5 65.9 56.9 52.3 70.9 72.9 54.0 54.9 54.7 54.9
arsenic-female-bladder 23.9 24.2 23.8 26.3 19.2 26.2 27.7 28.9 30.1 26.2 30.3 20.4 21.6 25.1 21.8 20.7 28.5 27.2 27.3 30.5 28.4 30.5 23.2 24.3 25.0 29.4 20.8 26.5 21.4 26.5
okcupid_stem 40.2 41.9 38.5 43.8 37.5 40.7 41.7 41.7 41.4 40.7 42.6 40.2 40.2 39.8 40.4 40.2 56.9 50.6 57.1 57.9 56.7 57.9 51.0 51.5 56.9 56.3 40.4 50.8 50.4 50.8
ecoli 63.6 56.9 63.0 54.3 58.9 65.9 67.2 67.8 67.5 63.6 66.7 68.2 64.5 64.0 66.4 64.8 88.3 87.9 88.8 88.9 89.0 88.9 72.7 69.7 87.9 86.3 68.6 71.0 71.8 71.0
pc4 30.0 33.9 32.8 28.4 26.2 34.7 39.6 38.3 39.5 35.4 38.7 34.3 37.1 35.2 35.3 35.7 50.6 46.7 40.4 43.9 43.4 43.9 35.0 39.2 38.3 45.0 33.4 34.8 36.7 34.8
bank-marketing 27.0 26.0 13.9 23.6 18.0 26.6 32.8 34.7 33.7 26.6 32.0 22.0 27.0 26.3 27.4 26.8 36.8 36.0 33.4 35.5 36.5 35.5 24.5 26.4 25.5 29.3 25.5 27.3 24.5 27.3
Diabetes-130-Hospitals_(Fairlearn) 11.4 11.8 11.3 12.6 11.3 11.6 11.7 12.0 11.8 11.5 11.7 11.5 11.5 11.6 11.5 11.5 13.3 11.7 13.4 13.0 12.9 13.0 11.5 11.5 11.2 11.3 11.4 11.5 11.5 11.5
Otto-Group-Product-Classification-Challenge 49.4 43.9 34.1 42.8 30.9 49.7 49.8 49.8 50.2 46.1 50.2 49.0 48.7 46.5 47.8 47.4 80.1 78.9 80.3 79.7 81.1 79.7 57.2 57.9 81.9 82.0 49.3 58.2 58.2 58.2
eucalyptus 63.2 49.7 38.5 41.9 39.5 61.1 59.4 57.8 59.4 61.1 59.5 61.2 62.5 62.4 62.7 61.5 85.1 82.4 84.1 84.8 86.2 84.8 66.6 66.0 87.4 87.0 60.9 69.1 65.6 69.1
pendigits 89.8 79.0 74.4 75.3 39.3 90.6 91.6 91.4 91.4 89.8 91.0 92.1 89.6 89.3 89.7 89.5 97.2 93.8 85.9 89.0 96.9 89.0 91.5 89.4 95.8 95.6 91.1 90.8 91.2 90.8
pc3 18.1 17.1 13.2 16.7 16.4 16.3 23.5 23.5 24.0 15.7 21.8 17.2 19.0 18.3 21.3 18.5 28.1 26.0 24.6 26.7 25.2 26.7 23.1 20.1 19.0 22.2 18.9 19.8 24.1 19.8
page-blocks-bin 69.2 57.5 45.5 57.6 30.9 71.1 73.7 73.4 72.2 71.5 73.2 67.2 69.5 69.2 72.3 65.9 78.6 78.7 67.1 72.0 78.5 72.0 76.3 70.0 78.0 78.7 69.3 75.6 73.3 75.6
optdigits 74.7 58.2 29.7 49.0 32.9 75.6 77.6 77.6 78.0 76.6 77.3 81.3 79.2 76.1 77.9 78.0 93.8 92.4 90.3 92.4 94.9 92.4 81.0 81.0 90.2 91.1 82.5 76.5 83.4 76.5
mfeat-karhunen 70.2 49.4 43.4 36.0 44.0 70.2 68.5 68.5 68.5 70.2 68.9 76.3 77.4 73.4 71.6 69.6 93.6 92.7 89.1 90.4 95.9 90.4 77.8 76.3 84.1 88.7 75.6 69.6 77.8 69.6
mfeat-fourier 90.8 81.0 43.6 71.6 40.4 90.8 90.8 90.8 90.8 90.8 90.8 93.5 93.5 95.3 97.2 97.2 99.1 97.8 99.1 98.7 99.6 98.7 94.4 94.8 97.3 98.7 96.3 92.7 94.4 92.7
mfeat-zernike 84.3 63.7 44.9 57.9 45.0 84.3 84.9 84.7 86.1 84.3 86.1 83.4 82.7 80.2 84.3 80.3 91.3 90.8 89.2 90.5 94.5 90.5 85.9 83.3 89.5 92.6 82.9 84.1 83.7 84.1
Pulsar-Dataset-HTRU2 70.7 45.3 36.3 53.0 18.9 71.0 72.0 71.9 71.6 71.0 70.4 69.2 62.3 70.7 70.2 63.2 80.3 78.4 70.5 74.1 76.1 74.1 69.2 59.7 79.7 79.7 69.4 71.1 69.7 71.1
vowel 86.2 54.5 37.4 67.1 39.8 86.2 86.3 86.2 86.3 88.8 88.2 78.4 84.7 91.8 83.2 82.0 94.9 82.8 78.0 80.7 96.9 80.7 82.3 86.5 97.0 96.9 78.8 85.4 81.5 85.4
[…]

### v2

@@ C.4 Comparison with BAF Benchmark
RUS $C$ IHT NM TL ENN RENN AKNN OSS NCR ROS SMT BSMT SSMT ASYN SENN STom SPE BC BRF $E$ UBS UBA OBS SMBS OBA SMBA AdaC AdaBS AsyBS CS XGB LGB CAT
bwin_amlb 36.4 34.4 35.9 36.9 35.6 35.3 39.6 39.6 36.4 35.4 38.6 39.5 36.3 37.1 35.4 35.9 36.0 35.9 38.6 37.2 40.7 39.0 38.0 39.0 38.3 38.7 36.6 37.6 36.2 34.7 36.2 34.9 37.0 36.9 36.2
mozilla4 82.9 78.3 76.7 65.6 66.9 82.0 81.7 80.9 80.1 81.9 81.7 83.4 81.8 82.1 82.1 79.5 77.1 81.2 90.3 89.5 88.5 89.5 89.6 89.5 84.1 81.7 90.4 89.7 83.7 83.8 83.7 83.4 88.6 89.1 88.7
mc2 40.7 43.9 38.4 39.8 43.9 43.7 46.9 45.9 41.2 43.7 45.9 42.1 45.9 43.8 50.4 50.6 39.6 40.0 50.9 46.4 47.0 51.8 51.1 51.8 54.2 55.1 51.9 52.2 48.2 46.7 48.2 41.1 44.0 43.7 43.6
wholesale-customers 72.5 70.0 68.0 60.2 68.1 71.6 76.5 76.6 75.4 72.2 76.5 72.9 74.0 73.3 77.6 71.6 75.5 70.2 83.6 81.3 81.6 82.6 81.7 82.6 71.6 77.8 83.1 82.3 74.2 75.9 74.2 74.4 81.4 80.5 80.3
vertebra-column 57.2 55.7 59.7 56.9 54.2 53.3 62.6 61.3 62.4 53.5 61.7 55.1 58.3 61.1 64.7 60.6 60.0 62.1 66.1 63.5 64.5 64.9 67.4 64.9 61.5 64.5 65.4 65.6 62.0 55.8 62.0 56.5 65.5 62.1 68.0
law-school-admission-bianry 37.4 37.5 34.3 40.0 37.0 38.2 39.5 40.4 40.2 38.2 40.0 37.1 37.6 37.6 37.7 37.7 41.3 37.7 42.7 41.6 43.1 44.1 42.9 44.1 39.2 39.4 42.7 42.1 39.4 39.0 39.4 37.4 43.6 44.2 44.6
bank32nh 46.4 45.4 44.0 47.0 45.7 45.7 47.9 48.0 48.2 46.7 47.8 45.4 45.7 46.4 47.4 46.0 45.5 45.9 57.2 57.4 55.7 56.6 56.4 56.6 46.8 45.7 54.9 56.4 47.2 46.3 47.2 45.7 57.6 58.0 58.0
elevators 56.9 52.0 46.0 47.4 45.4 56.2 56.6 56.5 55.3 56.4 56.6 55.6 56.6 56.6 56.2 56.8 56.0 56.3 71.1 70.3 65.8 67.4 68.3 67.4 56.6 57.7 69.6 70.0 57.2 56.0 57.2 55.9 74.8 73.6 73.3
cpu_small 70.9 69.1 69.5 57.8 61.5 72.1 73.3 72.7 72.5 72.7 72.9 72.0 71.7 70.4 71.3 69.5 71.8 72.4 80.6 79.4 78.3 79.0 80.0 79.0 71.4 72.1 80.2 79.9 71.6 71.3 71.6 71.0 81.2 81.6 81.2
Credit_Approval_Classification 38.1 36.3 36.6 39.4 38.6 41.2 41.2 39.2 39.7 41.2 40.3 37.6 38.8 38.1 39.1 40.0 37.5 38.0 46.6 44.7 46.1 47.5 46.0 47.5 39.0 38.7 45.4 46.4 39.0 40.6 39.0 39.1 45.4 46.1 46.3
house_8L 61.2 58.0 54.2 52.0 47.7 61.7 62.6 62.4 62.5 61.9 62.2 61.7 61.2 61.4 61.0 59.4 64.0 60.8 72.0 70.9 68.6 69.9 70.3 69.9 61.4 61.1 71.6 71.2 61.7 61.7 61.7 61.2 71.8 72.3 72.7
house_16H 61.3 58.0 55.6 51.6 46.7 61.6 62.1 62.3 62.2 61.6 62.1 60.1 60.2 61.3 59.9 58.8 61.6 60.0 72.5 71.6 69.3 70.6 70.7 70.6 59.2 60.7 72.1 71.4 61.6 59.1 61.6 59.3 72.2 73.6 73.2
phoneme 67.5 61.1 62.4 50.8 54.8 66.1 66.1 65.9 66.1 66.2 66.1 66.2 66.2 67.5 66.7 65.9 64.5 66.9 76.9 74.8 70.7 72.3 73.9 72.3 65.5 66.5 74.6 75.3 68.7 66.0 68.7 65.7 73.1 73.0 68.8
ilpd-numeric 33.8 35.4 38.6 38.4 30.8 33.6 40.2 39.6 39.4 33.8 41.4 38.4 38.3 37.2 34.8 35.8 38.2 36.8 43.6 38.2 45.0 44.0 43.8 44.0 33.5 37.2 35.0 39.2 32.6 36.5 32.6 38.2 36.3 37.5 32.4
planning-relax 27.7 31.7 27.4 28.5 30.9 31.0 29.9 30.3 29.8 30.1 31.4 27.6 33.0 31.1 30.7 33.3 27.9 31.4 28.8 28.1 28.5 30.1 34.0 30.1 29.4 32.3 31.9 32.1 29.4 30.4 29.4 30.7 31.6 28.9 28.6
MiniBooNE 71.0 69.2 65.0 58.2 61.3 71.7 70.4 68.5 69.5 71.6 70.9 70.6 70.5 69.3 70.3 69.6 70.3 70.6 82.6 81.5 78.9 80.1 77.6 80.1 77.7 77.5 81.0 81.3 77.5 77.6 77.5 70.6 83.5 82.8 83.1
machine_cpu 78.9 73.2 76.3 63.4 78.2 77.5 76.2 75.4 76.2 77.5 77.7 77.7 82.9 83.8 82.9 81.2 71.7 83.8 84.5 80.5 81.3 85.4 85.7 85.4 78.4 80.8 81.0 85.0 78.9 81.7 78.9 77.9 78.3 80.8 85.1
telco-customer-churn 36.3 38.0 36.0 39.3 29.0 38.4 41.8 43.3 42.4 38.5 41.5 37.5 38.2 38.9 38.6 39.0 44.0 38.8 45.7 44.7 46.0 46.7 46.0 46.7 40.9 40.0 45.3 45.9 37.9 39.0 37.9 37.5 44.4 46.1 46.0
haberman 30.3 28.1 27.6 32.6 32.2 34.4 36.5 34.8 34.2 34.7 36.5 30.6 32.2 32.2 30.3 30.6 33.1 29.4 33.6 31.9 35.1 36.0 32.7 36.0 32.3 34.2 33.1 36.4 30.5 31.7 30.5 31.0 28.2 29.2 31.2
vehicle 88.6 80.2 81.8 69.3 80.0 86.8 89.9 88.7 87.3 86.5 88.8 85.6 84.9 88.6 88.0 89.0 84.2 85.7 95.4 91.4 92.4 95.5 95.7 95.5 86.6 87.5 96.3 96.5 89.4 86.8 89.4 87.2 95.0 95.7 94.6
cpu 93.8 90.1 90.6 72.0 93.8 93.5 93.8 93.8 90.9 93.5 93.8 91.4 91.2 92.4 91.4 95.4 81.3 88.2 98.2 95.3 92.9 92.9 96.7 92.9 92.5 94.0 95.5 95.5 93.8 92.5 93.8 91.1 94.0 98.2 98.2
ada 45.0 45.1 43.4 41.1 36.1 46.9 50.6 51.5 50.4 47.3 49.0 46.5 46.8 45.1 45.4 44.8 49.5 46.8 57.4 56.8 54.6 57.1 55.4 57.1 46.8 46.8 58.0 57.7 45.9 46.3 45.9 45.6 58.3 58.8 58.8
adult 47.5 45.4 41.0 41.9 32.3 48.6 48.9 48.4 48.8 49.0 48.9 47.2 46.9 45.7 46.9 46.1 49.6 47.3 55.6 56.0 53.2 54.9 54.4 54.9 50.5 49.1 56.0 55.6 49.4 49.8 49.4 47.5 59.7 59.8 59.1
blood-transfusion-service-center 27.8 30.1 27.7 31.1 25.7 29.2 32.6 34.0 31.2 29.0 33.4 27.8 28.1 27.1 29.5 27.7 34.7 29.2 34.0 30.9 33.6 35.7 34.9 35.7 31.2 29.8 33.9 34.6 29.2 29.6 29.2 28.7 30.6 32.8 38.5
default-of-credit-card-clients 29.1 28.1 26.0 29.3 24.6 29.7 32.0 31.5 32.1 29.9 31.7 29.1 28.2 29.2 28.6 28.3 32.1 28.5 37.6 38.3 36.6 38.1 37.2 38.1 32.9 31.2 37.3 38.0 32.1 32.2 32.1 28.8 37.5 38.7 37.7
Customer_Churn_Classification 37.7 35.5 26.0 32.9 27.3 38.8 41.1 40.4 41.1 38.7 41.2 37.3 37.7 36.9 37.8 36.6 42.3 38.0 48.2 49.3 45.7 47.6 44.3 47.6 43.9 43.3 47.7 48.6 42.6 43.1 42.6 37.7 50.9 51.1 50.7
SPECTF 26.7 28.9 34.0 31.0 23.2 27.9 34.4 35.9 39.1 30.3 38.3 29.7 26.9 32.0 30.3 31.1 35.9 26.7 39.4 39.1 42.2 43.3 39.3 43.3 31.4 36.2 40.2 36.8 25.6 29.4 25.6 30.1 37.1 33.2 32.3
Medical-Appointment-No-Shows 25.0 25.6 20.1 27.2 23.2 25.1 26.4 27.3 26.8 25.1 26.3 24.9 24.3 24.6 24.6 24.3 26.6 24.6 28.2 26.8 28.8 28.3 28.0 28.3 24.7 24.4 25.0 25.8 25.0 24.8 25.0 24.8 21.9 20.5 21.5
JapaneseVowels 85.6 75.4 76.6 64.6 66.3 85.6 85.9 86.1 85.8 85.5 85.6 85.4 85.9 85.6 84.1 83.9 83.8 84.7 94.5 91.8 88.5 91.7 93.9 91.7 86.5 86.0 92.4 93.3 86.2 86.5 86.2 86.3 95.0 95.2 92.9
ibm-employee-attrition 24.2 21.1 18.8 23.8 18.2 22.2 25.1 25.1 24.8 22.4 24.8 21.8 25.8 25.3 24.6 24.4 23.4 20.8 36.6 33.9 32.3 34.9 34.4 34.9 22.8 24.6 26.1 29.9 23.5 22.4 23.5 20.7 29.8 30.4 36.4
first-order-theorem-proving 30.8 26.3 24.2 26.4 22.4 29.6 28.1 28.1 27.9 29.7 28.1 29.8 30.5 30.8 30.9 30.2 27.5 29.6 51.3 48.6 51.2 51.7 51.8 51.7 51.2 50.6 53.0 52.9 51.2 51.0 51.2 29.9 52.9 52.6 49.4
user-knowledge 71.1 63.4 62.6 54.8 60.1 67.5 69.0 69.0 67.2 67.5 69.0 68.5 74.2 71.1 67.3 71.1 64.5 74.6 89.4 89.0 86.0 86.2 90.2 86.2 72.4 73.9 86.7 86.7 74.0 74.2 74.0 69.2 87.8 87.2 88.2
online-shoppers-intention 38.1 36.7 24.6 31.3 25.2 38.9 41.8 40.9 41.3 39.1 41.3 38.6 38.7 37.9 40.0 39.2 41.1 39.4 50.0 50.4 46.2 48.1 47.7 48.1 37.3 39.9 47.5 50.5 37.2 37.6 37.2 37.7 47.2 49.1 48.4
kc1 24.2 24.4 14.9 23.9 20.7 24.3 25.9 27.5 23.7 24.5 26.7 23.4 27.8 26.9 29.0 26.6 25.4 26.4 31.0 26.8 28.3 30.2 30.8 30.2 29.4 28.6 28.3 30.6 25.2 24.9 25.2 22.1 29.7 29.5 28.8
thoracic-surgery 17.1 14.9 16.8 17.6 17.8 17.7 17.8 18.1 17.4 17.6 20.4 18.0 19.0 19.2 20.3 19.7 16.1 19.7 20.5 20.7 20.3 20.0 22.5 20.0 16.7 20.6 16.1 16.6 16.9 17.4 16.9 16.9 16.4 16.9 14.9
UCI_churn 53.4 35.3 21.4 27.2 31.7 52.2 54.4 54.5 54.4 52.6 54.2 54.2 53.0 53.5 52.2 51.4 45.2 53.0 74.6 71.0 58.3 65.9 65.5 65.9 56.9 52.3 70.9 72.9 54.9 54.7 54.9 54.0 73.8 73.6 74.2
arsenic-female-bladder 23.9 24.2 23.8 26.3 19.2 26.2 27.7 28.9 30.1 26.2 30.3 20.4 21.6 25.1 21.8 20.7 26.3 21.8 28.5 27.2 27.3 30.5 28.4 30.5 23.2 24.3 25.0 29.4 26.5 21.4 26.5 20.8 26.1 26.8 22.7
okcupid_stem 40.2 41.9 38.5 43.8 37.5 40.7 41.7 41.7 41.4 40.7 42.6 40.2 40.2 39.8 40.4 40.2 43.6 40.6 56.9 50.6 57.1 57.9 56.7 57.9 51.0 51.5 56.9 56.3 50.8 50.4 50.8 40.4 59.4 60.1 59.7
ecoli 63.6 56.9 63.0 54.3 58.9 65.9 67.2 67.8 67.5 63.6 66.7 68.2 64.5 64.0 66.4 64.8 62.1 64.1 88.3 87.9 88.8 88.9 89.0 88.9 72.7 69.7 87.9 86.3 71.0 71.8 71.0 68.6 86.9 88.6 89.0
pc4 30.0 33.9 32.8 28.4 26.2 34.7 39.6 38.3 39.5 35.4 38.7 34.3 37.1 35.2 35.3 35.7 38.2 34.2 50.6 46.7 40.4 43.9 43.4 43.9 35.0 39.2 38.3 45.0 34.8 36.7 34.8 33.4 40.0 41.0 41.7
bank-marketing 27.0 26.0 13.9 23.6 18.0 26.6 32.8 34.7 33.7 26.6 32.0 22.0 27.0 26.3 27.4 26.8 33.3 27.5 36.8 36.0 33.4 35.5 36.5 35.5 24.5 26.4 25.5 29.3 27.3 24.5 27.3 25.5 28.4 30.7 28.8
Diabetes-130-Hospitals_(Fairlearn) 11.4 11.8 11.3 12.6 11.3 11.6 11.7 12.0 11.8 11.5 11.7 11.5 11.5 11.6 11.5 11.5 11.9 11.5 13.3 11.7 13.4 13.0 12.9 13.0 11.5 11.5 11.2 11.3 11.5 11.5 11.5 11.4 11.2 11.2 11.3
Otto-Group-Product-Classification-Challenge 49.4 43.9 34.1 42.8 30.9 49.7 49.8 49.8 50.2 46.1 50.2 49.0 48.7 46.5 47.8 47.4 48.3 48.7 80.1 78.9 80.3 79.7 81.1 79.7 57.2 57.9 81.9 82.0 58.2 58.2 58.2 49.3 83.6 83.4 79.2
eucalyptus 63.2 49.7 38.5 41.9 39.5 61.1 59.4 57.8 59.4 61.1 59.5 61.2 62.5 62.4 62.7 61.5 56.9 62.1 85.1 82.4 84.1 84.8 86.2 84.8 66.6 66.0 87.4 87.0 69.1 65.6 69.1 60.9 87.9 87.2 86.1
pendigits 89.8 79.0 74.4 75.3 39.3 90.6 91.6 91.4 91.4 89.8 91.0 92.1 89.6 89.3 89.7 89.5 89.9 89.6 97.2 93.8 85.9 89.0 96.9 89.0 91.5 89.4 95.8 95.6 90.8 91.2 90.8 91.1 97.8 97.4 97.4
pc3 18.1 17.1 13.2 16.7 16.4 16.3 23.5 23.5 24.0 15.7 21.8 17.2 19.0 18.3 21.3 18.5 20.9 17.0 28.1 26.0 24.6 26.7 25.2 26.7 23.1 20.1 19.0 22.2 19.8 24.1 19.8 18.9 21.7 20.6 19.8
page-blocks-bin 69.2 57.5 45.5 57.6 30.9 71.1 73.7 73.4 72.2 71.5 73.2 67.2 69.5 69.2 72.3 65.9 69.5 68.4 78.6 78.7 67.1 72.0 78.5 72.0 76.3 70.0 78.0 78.7 75.6 73.3 75.6 69.3 77.2 76.4 78.0
optdigits 74.7 58.2 29.7 49.0 32.9 75.6 77.6 77.6 78.0 76.6 77.3 81.3 79.2 76.1 77.9 78.0 78.6 78.7 93.8 92.4 90.3 92.4 94.9 92.4 81.0 81.0 90.2 91.1 76.5 83.4 76.5 82.5 93.8 94.1 94.1
mfeat-karhunen 70.2 49.4 43.4 36.0 44.0 70.2 68.5 68.5 68.5 70.2 68.9 76.3 77.4 73.4 71.6 69.6 75.2 73.2 93.6 92.7 89.1 90.4 95.9 90.4 77.8 76.3 84.1 88.7 69.6 77.8 69.6 75.6 91.3 93.7 90.9
mfeat-fourier 90.8 81.0 43.6 71.6 40.4 90.8 90.8 90.8 90.8 90.8 90.8 93.5 93.5 95.3 97.2 97.2 91.5 91.0 99.1 97.8 99.1 98.7 99.6 98.7 94.4 94.8 97.3 98.7 92.7 94.4 92.7 96.3 97.8 98.7 98.7
mfeat-zernike 84.3 63.7 44.9 57.9 45.0 84.3 84.9 84.7 86.1 84.3 86.1 83.4 82.7 80.2 84.3 80.3 82.5 78.3 91.3 90.8 89.2 90.5 94.5 90.5 85.9 83.3 89.5 92.6 84.1 83.7 84.1 82.9 90.9 91.8 90.0
Pulsar-Dataset-HTRU2 70.7 45.3 36.3 53.0 18.9 71.0 72.0 71.9 71.6 71.0 70.4 69.2 62.3 70.7 70.2 63.2 62.5 62.1 80.3 78.4 70.5 74.1 76.1 74.1 69.2 59.7 79.7 79.7 71.1 69.7 71.1 69.4 80.0 80.0 80.2
vowel 86.2 54.5 37.4 67.1 39.8 86.2 86.3 86.2 86.3 88.8 88.2 78.4 84.7 91.8 83.2 82.0 84.7 84.7 94.9 82.8 78.0 80.7 96.9 80.7 82.3 86.5 97.0 96.9 85.4 81.5 85.4 78.8 97.0 92.9 96.0
[…]